I need to find out where the equations x + 2y=1and -2x+ y=-4 intersect. I have tried and my lines are not connecting!!
Answer by Alan3354(69443)

I need to find out where the equations x + 2y=1 and -2x+ y=-4 intersect. I have tried and my lines are not connecting!!
---------------------
x + 2y=1
-2x+ y=-4
-------
2x + 4y = 2 (1st eqn times 2)
-2x+ y = -4
---------------- Add
5y = -2
y = -2/5
---------------
x + 2y = 1
x + (-4/5) = 1
x = 9/5
--------------
they intersect at the point (9/5,-2/5)